# Adversarial Score Identity Distillation: Rapidly Surpassing the Teacher in One Step

### v2

@@ 6 Conclusion
In this paper, we present Score identity Distillation with Adversarial Loss (SiDA), an innovative framework that optimizes diffusion model distillation by integrating the strengths of Score identity Distillation (SiD) and Diffusion-GAN. SiDA accelerates the generation process, significantly reducing the number of iterations required compared to conventional score distillation methods, while producing high-quality images. Extensive tests across various datasets—CIFAR-10, ImageNet (64x64 and 512x512), FFHQ, AFHQ-v2—and different scales of pretrained EDM2 models demonstrate that SiDA consistently surpasses both the base teacher diffusion models and other distilled variants. This is validated by superior Fréchet Inception Distance (FID) scores and other performance metrics. Notably, SiDA achieves these results without utilizing classifier-free guidance, suggesting potential for further enhancement with its integration. Specifically, incorporating the long-and-short guidance strategy as described by Zhou et al. (2024a) could further refine SiDA’s performance. Future work will explore expanding SiDA’s methodology to additional generative tasks and refining its capabilities through innovative distillation techniques.
@@
Appendix A Algorithm Box
Algorithm 1 Adversarial Score identity Distillation (SiDA)
1:  Input: Pretrained score network $f_{\phi}$, generator $G_{\theta}$, generator score network $f_{\psi}$, $\sigma_{\text{init}}=2.5$, $t_{\max}=800$, $\alpha=1.2$, $\lambda_{\text{sid}}=100$, $\lambda_{\text{adv}}=0.01$, image size $(W,H,C_{c})$, latent discriminator map size $(W^{\prime},H^{\prime})$.
3:  repeat
6:     Update $\psi$ with 9:
7:         $\mathcal{L}_{\psi}^{(\text{sida})}={\gamma(t)}(\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}+{L}_{\psi}^{(\text{adv})})$
8:         $\psi=\psi-\eta\nabla_{\psi}\mathcal{L}_{\psi}^{(\text{sida})}$
10:     if num_imgs $\geq$ 100K then
11:        Sample $t\sim\mbox{Unif}[0,t_{\max}/1000]$, compute $\sigma_{t}$, $\omega_{t}$ and $a_{t}$ as defined in Zhou et al. (2024b)
12:        Set $b=0$ if num_imgs $\leq$200k and $b=1$ otherwise
13:        Update $G_{\theta}$ with 8:
15:                  $\left.+\frac{\omega(t)a_{t}^{2}}{{\sigma_{t}^{4}}}(f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t))^{T}(f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})\right)+\frac{b}{2}\lambda_{\text{adv}}\frac{\omega(t)a_{t}^{2}}{2\sigma_{t}^{4}}{C_{c}WH}\mathcal{L}_{\theta}^{(\text{adv})}$
17:     end if
18:  until the FID plateaus or the training budget is exhausted
19:  Output: $G_{\theta}$
Algorithm 2 SiD2A: SiD initilized SiDA
1:  Extra Input: Pretrained SiD generator $\theta_{\text{sid}}$
3:  repeat
4:     The same as SiDA
5:  until the FID plateaus or the training budget is exhausted
6:  Output: $G_{\theta}$
Appendix B Additional Training and Evaluation Details
Table 5: Analogous to Table 2 for FFHQ 64x64.
Table 6: Analogous to Table 2 for AFHQ-v2 64x64.
[…]
Method CFG NFE XS (125M) S (280M) M (498M) L ( [^7] M) XL (1.1B) XXL (1.5B)
EDM2 N 63 103.39 68.64 58.44 52.25 45.96 42.84
EDM2 Y 63 79.94 52.32 41.98 38.20 35.67 33.09
SiD N 1 91.75 $\pm$ 0.55 65.08 $\pm$ 0.32 55.92 $\pm$ 0.25 56.25 $\pm$ 0.40 55.87 $\pm$ 0.21 -
SiDA N 1 93.67 $\pm$ 0.40 68.76 $\pm$ 0.45 53.40 $\pm$ 0.60 48.47 $\pm$ 0.45 46.47 $\pm$ 0.27 -
SiD2A N 1 86.96 $\pm$ 0.32 62.19 $\pm$ 0.36 49.01 $\pm$ 0.54 45.92 $\pm$ 0.33 - -
The hyperparameters customized for our study are outlined in Table 8, with all remaining settings consistent with those in the EDM code (Karras et al., 2022) and SiD code (Zhou et al., 2024b).
@@ Appendix B Additional Training and Evaluation Details
Table 8: Hyperparameter settings and comparison of distillation time, and memory usage between SiD and SiDA, using 8 NVIDIA V100 GPUs (16 GB each) or 8 NVIDIA H100 GPUs (80 GB each).
@@
Loss scaling of ($\lambda_{\text{sid}},\,\lambda_{\text{adv}}$) (100, 0.01)

### v4

@@ 6 Conclusion
In this paper, we present Score identity Distillation with Adversarial Loss (SiDA), an innovative framework that optimizes diffusion model distillation by integrating the strengths of Score identity Distillation (SiD) and Diffusion-GAN. SiDA accelerates the generation process, significantly reducing the number of iterations required compared to conventional score distillation methods, while producing high-quality images. Extensive tests across various datasets—CIFAR-10, ImageNet (64x64 and 512x512), FFHQ, AFHQ-v2—and different scales of pretrained EDM2 models demonstrate that SiDA consistently surpasses both the base teacher diffusion models and other distilled variants. This is validated by superior Fréchet Inception Distance scores and other performance metrics. Notably, SiDA achieves these results without utilizing classifier-free guidance, suggesting potential for further enhancement with its integration. Specifically, incorporating the long-and-short guidance strategy as described by Zhou et al. (2024a) could further refine SiDA’s performance. Future work will explore expanding SiDA’s methodology to additional generative tasks and refining its capabilities through innovative distillation techniques.
Acknowledgments
M. Zhou, H. Zheng, Y. Gu, and Z. Wang acknowledge the support of NSF-IIS 2212418 and NIH-R37 CA271186.
@@ Acknowledgments
Appendix A Experimetal Settings
Implementation Details. We implement SiDA using the SiD codebase (Zhou et al., 2024b), incorporating essential functions adapted from both EDM (Karras et al., 2022) and EDM2 (Karras et al., 2024). Detailed adaptations of EDM2 modifications within SiD and SiDA are provided in Appendix B.2, specifically addressing our handling of forced weight normalization and data-specific weighting for the loss terms introduced by EDM2 to enhance EDM. We did not incorporate classifier-free guidance (Ho & Salimans, 2022), use two teacher models, or set exponential moving average (EMA) parameters post-hoc, all of which are important for EDM2’s performance improvements over EDM. Even without these techniques, SiD already performs on par with EDM2 teacher models, while SiDA and SiD2A significantly outperform the teacher. These additional techniques could potentially further enhance SiD and SiDA, which we leave for future work.
The score estimation network, $f_{\psi}$, is initialized using the same architecture and parameters as the pre-trained teacher score network, $f_{\phi}$, from either EDM or EDM2. For the initialization of the student generator, $G_{\theta}$, we explore two settings: one where $G_{\theta}$ is initialized with the same parameters as the pre-trained EDM or EDM2 teacher score network, and another where it is initialized using the distilled generator from SiD, referred to as SiD-SiDA (SiD2A). The hyperparameters customized for our study are outlined in Table 7, with all remaining settings consistent with those in the EDM code (Karras et al., 2022), EDM2 code (Karras et al., 2024), and SiD code (Zhou et al., 2024b).
Our training process unfolds in three stages: 1) For the first 100k images, we exclusively train $f_{\psi}$ to stabilize the score estimation. 2) For the subsequent 100k images, both $f_{\psi}$ and $G_{\theta}$ are trained using the SiD generator loss (6), integrating the distilled insights. 3) For all remaining images, we update $f_{\psi}$ and $G_{\theta}$ with the SiDA generator loss (8), which incorporates adversarial adjustments to refine the generation process. Throughout the training, the score estimation network $f_{\psi}$ is consistently trained with the SiDA fake score loss (9), ensuring that the adversarial components are properly integrated.
We observe that using only the adversarial loss, without the SiD component in the combined loss, leads to rapid divergence, even when the single-step generator is initialized from SiD. This observation suggests that while the added adversarial components cannot stand alone, they effectively complement the SiD loss, which operates in a data-free manner, allowing it to leverage available training data to correct the teacher’s bias and thus enhance the distillation speed and performance.
Parameter Settings. To balance the SiD and adversarial components in both the generator and fake-score losses, we set $\lambda_{\text{sid}},\lambda_{\text{adv},\theta},\lambda_{\text{adv},\psi}$ to align the scales of the four different losses. This keeps them approximately within the range of 1 to 10,000 to prevent fp16 overflow/underflow. The settings in Tables 6 and 7 achieve this. For EDM distillation, we reused the SiD parameters. For EDM2, we matched the batch size to the teacher’s and adapted $\lambda_{\text{adv},\psi}$ from 1 to 100 to roughly equate the adversarial loss scale with the diffusion distillation loss for the fake score network. Initially, we used a learning rate of 4e-6 but switched to 5e-5 for faster convergence in EDM2-XS. Our experience indicates that SiD and SiDA are not overly sensitive to these settings as long as loss scales are comparable, suggesting robustness. While fine-tuning these hyper-parameters could potentially further improve outcomes, our focus wasn’t on hyperparameter optimization given the already achieved state-of-the-art performance across various settings.
@@ Appendix A Experimetal Settings
Table 6: Hyperparameter settings of SiDA and comparison of distillation time, and memory usage between SiD and SiDA, using 8 NVIDIA V100 GPUs (16 GB each) or 8 NVIDIA H100 GPUs (80 GB each).
@@
Loss scaling of ($\lambda_{\text{sid}},\,\lambda_{\text{adv},\theta},\,\lambda_{\text{adv},\psi}$) (100, 0.01, 1)
@@
Table 7: Hyperparameter settings and comparison of distillation time, and memory usage between SiD and SiDA on distilling EDM2 pretrained on ImageNet 512x512, using 16 NVIDIA A100 GPUs (40 GB each) or 8 NVIDIA H100 GPUs (80 GB each).
Method Hyperparameters EDM2-XS EDM2-S EDM2-M EDM2-L EDM2-XL EDM2-XXL
Batch size 2048
Batch size per GPU 64 64 32 32 16 2
GPUs 16xA100 8xH100 8xH100 8xH100 8xH100 8xH100
Gradient accumulation round 2 4 8 8 16 128
Learning rate of ($\psi$, $\theta$) 5e-5
Loss scaling of ($\lambda_{\text{sid}},\,\lambda_{\text{adv},\theta},\,\lambda_{\text{adv},\psi}$) (100, 0.01, 100)
ema 2
fp16 True
Optimizer Adam ($\beta_{1}$, $\beta_{2}$, eps) (0, 0.999, 1e-6)
$\alpha$ $1.0$
$\sigma_{\text{init}}$ 2.5
$t_{\max}$ 800
dropout The same as in EDM2 for each corresponding model size
SiD max memory allocated per GPU 31.5 51.3 49.4 68.8 72.2 74.6
max memory in GB reserved per GPU 32.3 52.4 50.0 70.0 73.4 75.1
[…]
SiDA max memory allocated per GPU 31.5 51.3 49.4 68.9 72.2 74.6
max memory in GB reserved per GPU 34.2 52.4 52.0 70.1 73.4 75.1
[…]
Datasets and Teacher Models. To comprehensively evaluate the effectiveness of SiDA, we first utilize four representative datasets of varying scales and resolutions and the EDM diffusion models pretrained on them, as discussed in Karras et al. (2022). These datasets include CIFAR-10 ($32\times 32$; both conditional and unconditional) (Krizhevsky et al., 2009), ImageNet 64x64 (Deng et al., 2009), FFHQ ($64\times 64$) (Karras et al., 2019), and AFHQ-v2 ($64\times 64$) (Choi et al., 2020).
Additionally, we distill EDM2 diffusion models of varying sizes that were pretrained on ImageNet 512x512, following details provided by Karras et al. (2024). We compare these distilled models to their corresponding teacher models and to a concurrent work of Lu & Song (2024) on EDM2 distillation. The models range from 125M million to 1.5 billion parameters. By utilizing these diverse datasets and model scales, we thoroughly assess the performance of SiD, SiDA, and SiD2A across different content types and complexities, ensuring a robust evaluation of their generative capabilities.
Evaluation Protocol. We assess the quality of image generation using the Fréchet Inception Distance (FID) and Inception Score (IS) (Salimans et al., 2016). Following Karras et al. (2024), we also evaluate models trained on ImageNet 512x512 using $\text{FD}_{\text{DINOv2}}$, the Fréchet distance computed in the feature space of DINOv2 (Oquab et al., 2023). Consistent with the methodology of Karras et al. (2019; 2022; 2024), these scores are calculated using 50,000 generated samples, with the training dataset used by the EDM or EDM2 teacher model [^3] https://github.com/NVlabs/edm serving as the reference. Additionally, we evaluate SiDA on ImageNet 64x64 using the Precision and Recall metrics (Kynkäänniemi et al., 2019), where both metrics are calculated using a predefined reference batch [^4] https://openaipublic.blob.core.windows.net/diffusion/jul-2021/ref_batches/imagenet/64/VIRTUAL_imagenet64_labeled.npz [^5] https://github.com/openai/guided-diffusion/tree/main/evaluations (Dhariwal & Nichol, 2021a; Nichol & Dhariwal, 2021; Song et al., 2023; Song & Dhariwal, 2023).
In line with SiD, we periodically evaluate the FID during distillation and retain the generators with the lowest FID. To ensure accuracy and statistically robust comparisons, we perform re-evaluations across 10 independent runs to obtain more reliable metrics. We recommend this rigorous approach over reporting only the best metrics observed across individual runs or during the distillation process, as the latter can result in biased metrics that are consistently lower than the mean.
Appendix B Algorithm Details
B.1 Algorithm Box
Algorithm 1 Adversarial Score identity Distillation (SiDA)
1:  Input: Pretrained score network $f_{\phi}$, generator $G_{\theta}$, generator score network $f_{\psi}$, $\sigma_{\text{init}}=2.5$, $t_{\max}=800$, $\alpha=1.2$, $\lambda_{\text{sid}}=100$, $\lambda_{\text{adv},\theta}=0.01$, $\lambda_{\text{adv},\psi}=1$, image size $(W,H,C_{c})$, latent discriminator map size $(W^{\prime},H^{\prime})$.
3:  repeat
6:     Update $\psi$ with 9:
7:         $\mathcal{L}_{\psi}^{(\text{sida})}={\gamma(t)}(\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{\text{adv},\psi}}{L}_{\psi}^{(\text{adv})})$
8:         $\psi=\psi-\eta\nabla_{\psi}\mathcal{L}_{\psi}^{(\text{sida})}$
10:     if num_imgs $\geq$ 100K then
11:        Sample $t\sim\mbox{Unif}[0,t_{\max}/1000]$, compute $\sigma_{t}$, $\omega_{t}$, and $a_{t}$ as defined in Zhou et al. (2024b)
12:        Set $b=0$ if num_imgs $\leq$200k and $b=1$ otherwise
13:        Update $G_{\theta}$ with 8:
15:                  $\left.+\frac{\omega(t)a_{t}^{2}}{{\sigma_{t}^{4}}}(f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t))^{T}(f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})\right)+\frac{b}{2}\lambda_{\text{adv},\theta}\frac{\omega(t)a_{t}^{2}}{2\sigma_{t}^{4}}{C_{c}WH}\mathcal{L}_{\theta}^{(\text{adv})}$
17:     end if
18:  until the FID plateaus or the training budget is exhausted
19:  Output: $G_{\theta}$
Algorithm 2 SiDA for EDM2
1:  Extra Input: $\text{LogVar}_{\text{Flag}}\in\{\text{True},\text{False}\}$$\text{ForceNorm}_{\text{Flag}}\in\{\text{True},\text{False}\}$, $\lambda_{\text{adv},\psi}=100$
2:  repeat
3:     The same as SiDA for EDM with the following modifications:
4:     if $\text{LogVar}_{\text{Flag}}$ is True then
5:        Modify the fake score network loss as $\mathcal{L}_{\psi}^{(\text{sida})}=\frac{{\gamma(t)}}{e^{\text{logvar}}}(\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}+\text{logvar}+\lambda_{\text{adv},\psi}{L}_{\psi}^{(\text{adv})})$
6:     end if
7:     if $\text{ForceNorm}_{\text{Flag}}$ is True then
8:        During training, apply forced weight normalization used by EDM2 for SiD, and a modified version of it to prevent gradient backpropagation error for SiDA
9:     end if
10:  until the FID plateaus or the training budget is exhausted
11:  Output: $G_{\theta}$
Algorithm 3 SiD2A: SiD initilized SiDA (applicable for both EDM and EDM2)
1:  Extra Input: Pretrained SiD generator $\theta_{\text{sid}}$
3:  repeat
4:     The same as SiDA
5:  until the FID plateaus or the training budget is exhausted
6:  Output: $G_{\theta}$
B.2 Forced Weight Normalization with Pre-Hook
An important training technique used by EDM2 to enhance EDM is the application of forced weight normalization, where each weight vector is explicitly normalized to unit variance before every training step. This method is found to be fully compatible with the training of both the generator and fake score network in SiD and does not introduce gradient backpropagation errors. However, in SiDA, where adversarial loss is incorporated, we find that this approach often leads to computational errors during gradient backpropagation.
We have traced the source of the problem to the MPConv class of the EDM2 code, which modifies self.weight in-place during the forward pass. A possible reason is that even though torch.no_grad() was used, in-place operations on parameters that require gradients can interfere with PyTorch’s autograd system, leading to errors during backpropagation.
To address these issues, we provide two options:
  1. 1.
    Disable Forced Weight Normalization: Completely remove the in-place normalization step from the training process. This avoids the gradient backpropagation errors but also forgoes the potential benefits of forced weight normalization.
  2. 2.
    Use a Pre-Forward Hook to Avoid In-Place Operations: Instead of modifying self.weight in-place during the forward pass, we utilize a pre-forward hook to normalize the weights before each forward pass. Executed before the forward method of a module, a pre-hook in PyTorch allows for consistent normalization of the weights without interfering with the autograd system. Employing this strategy enables forced weight normalization without causing gradient backpropagation problems in SiDA.
By implementing the second option, we maintain the potential advantages of forced weight normalization while ensuring stable training with adversarial loss in SiDA.
In our experiments, we assessed two options during the distillation of EDM2-XS, EDM2-S, and EDM2-M and observed no significant performance differences. For SiD, we retained the original forced weight normalization. For SiDA, we generally opted for forced weight normalization with pre-hook (option 2), except for EDM2-S and EDM2-XXL where we used no forced weight normalization (option 1). Similarly, SiD2A defaults to option 2, but switches to option 1 for EDM2-S. These variations in choices are not intended to maximize performance for each setting, but rather to illustrate that the choice of normalization may not significantly impact results, or at least not impede their ability to outperform the teacher.
Another clear difference between EDM2 and EDM is the use of a one-layer MLP to produce the weights for each data-specific loss term. This MLP is trained alongside the main denoising network and discarded afterward. In our experiments, we did not observe a clear difference in distillation performance with or without this MLP. Therefore, we enable it by default when forced weight normalization is applied and disable it when no forced normalization is used.
Below, we provide the code for the original and modified versions of the MPConv class.
The Original MPConv Class of EDM2 That Includes Forced Weight Normalization
@persistence.persistent_class
class MPConv(torch.nn.Module):
def __init__(self, in_channels, out_channels, kernel):
super().__init__()
self.out_channels = out_channels
self.weight = torch.nn.Parameter(torch.randn(out_channels, in_channels, *kernel))
def forward(self, x, gain=1):
w = self.weight.to(torch.float32)
if self.training:
with torch.no_grad():
self.weight.copy_(normalize(w)) # forced weight normalization
w = normalize(w) # traditional weight normalization
w = w * (gain / np.sqrt(w[0].numel())) # magnitude-preserving scaling
w = w.to(x.dtype)
if w.ndim == 2:
return x @ w.t()
assert w.ndim == 4
return torch.nn.functional.conv2d(x, w, padding=(w.shape[-1]//2,))
Listing 1: Original MPConv class with forced weight normalization
The Modified MPConv Class Introducing an Optional Modified Forced Weight Normalization When Adversarial Loss Is Applied
@persistence.persistent_class
class MPConv(torch.nn.Module):
def __init__(self, in_channels, out_channels, kernel,force_normalization,use_gan):
super().__init__()
self.out_channels = out_channels
self.weight = torch.nn.Parameter(torch.randn(out_channels, in_channels, *kernel))
self.force_normalization =force_normalization
self.use_gan = use_gan
# Register the forward pre-hook
if self.use_gan and self.force_normalization:
self.register_forward_pre_hook(self._apply_forced_weight_normalization)
def _apply_forced_weight_normalization(self, module, input):
# Only apply during training
if self.training:
with torch.no_grad():
w = self.weight.to(torch.float32)
w_normalized = normalize(w)
self.weight.copy_(w_normalized)
def forward(self, x, gain=1):
w = self.weight.to(torch.float32)
if self.training and not self.use_gan:
with torch.no_grad():
self.weight.copy_(normalize(w)) # forced weight normalization
w = normalize(w) # Traditional weight normalization
w = w * (gain / np.sqrt(w[0].numel())) # Magnitude-preserving scaling
w = w.to(x.dtype)
if w.ndim == 2:
return x @ w.t()
assert w.ndim == 4
return torch.nn.functional.conv2d(x, w, padding=(w.shape[-1] // 2))
Listing 2: Modified MPConv class with pre-forward hook
Appendix C Additional Training and Evaluation Details
Table 8: Analogous to Table 2 for FFHQ 64x64.
Table 9: Analogous to Table 2 for AFHQ-v2 64x64.
Figure 7: Evolution of FID scores for the SiDA generator during the distillation of the EDM2-XS teacher model, pretrained on ImageNet 512×512. Training utilized either 5% or 100% of the real training images, with a batch size of 2048 and $\alpha=1.0$. Horizontal lines indicate the performance of EDM2 without classifier-free guidance (CFG) in purple and with CFG (using 63 NFEs) in red.
[…]
Method CFG NFE XS (125M) S (280M) M (498M) L ( [^7] M) XL (1.1B) XXL (1.5B)
EDM2 N 63 103.39 68.64 58.44 52.25 45.96 42.84
EDM2 Y 63x2 79.94 52.32 41.98 38.20 35.67 33.09
SiD N 1 91.75 $\pm$ 0.55 65.08 $\pm$ 0.32 55.92 $\pm$ 0.25 56.25 $\pm$ 0.40 52.47 $\pm$ 0.39 56.15 $\pm$ 0.41
SiDA N 1 93.67 $\pm$ 0.40 68.76 $\pm$ 0.45 53.40 $\pm$ 0.60 48.47 $\pm$ 0.45 44.47 $\pm$ 0.39 50.22 $\pm$ 0.43
SiD2A N 1 86.96 $\pm$ 0.32 62.19 $\pm$ 0.36 49.01 $\pm$ 0.54 46.80 $\pm$ 0.29 43.89 $\pm$ 0.51 44.52 $\pm$ 0.37
Figure 8: The SiD method highlights rapid advancements in distilling EDM2-XL pretrained on ImageNet 512x512, utilizing a batch size of 2048 and a learning rate of 5e-5. Illustrations are provided using random generations with class labels: 0 for tench and 1 for goldfish. The series of images, generated from the same set of random noises post-training the SiDA generator with varying counts of synthesized images, illustrates progressions at 0, 0.2, 0.5, 1, 2, 4,10, 20, 100, and 200 million images. These are equivalent to 0, 100, 250, 500, 1k, 2K, 5K, 10k, 50k, and 100k training iterations respectively, organized from the top left to the bottom right. The blue doshed curve in Panel (e) of Figure 5 details the progression of FIDs for SiD on EDM2-XL.
Figure 9: Analogous plot to Figure 8 for the the newly proposed SiDA method, with two more classes introduced for visulization: 2 for great white shark, and 3 for tiger shark. The black dotted curve in Panel (e) of Figure 5 details the progression of FIDs for SiDA on EDM2-XL.
Figure 10: Analogous plot to Figure 10 for the the newly proposed SiD2A method. The series of images, generated from the same set of random noises post-training the SiDA generator with varying counts of synthesized images, illustrates progressions at 0, 4, 10, 20, 40 million images. These are equivalent to 0, 2K, 5K, 10k, 20k training iterations respectively, organized from the left to right. The orange solid curve in Panel (e) of Figure 5 details the progression of FIDs for SiD2A on EDM2-XL.